OpenStudy (anonymous):

In humans, there is a phenotype for curly hair and a phenotype for straight hair. The homozygous dominant genotype CC is for curly hair, and the homozygous recessive genotype cc is for straight hair. Heterozygous Cc is a phenotype for wavy hair. In a hypothetical population, 36% of people are born with straight hair (cc). What percentage of this population will have wavy hair?

OpenStudy (anonymous):

Using the Hardy Weinberg equation p^2+2pq+q^2=1 and p+q=1 q^2 = .36 (the proportion of homozygous recessive) q = .6 p = .4 2pq = .48 48% of the population will have wavy hair.